# Provably Finding a Hidden Dense Submatrix among Many Planted Dense Submatrices via Convex Programming

### v1

@@ 3.1 Optimality Conditions
$\mathrm{Tr}(\boldsymbol{\Lambda}^{T}(\boldsymbol{\bar{X}}-\boldsymbol{1}\boldsymbol{1}^{T}))$ $=0$ (8b)
$\mathrm{Tr}(\boldsymbol{\Xi}^{T}\boldsymbol{\bar{Y}})$ $=0$ (8c)
$\boldsymbol{W}^{T}\boldsymbol{\bar{u}}=\boldsymbol{0},\quad\boldsymbol{W}\boldsymbol{\bar{v}}=\boldsymbol{0},\quad\|\boldsymbol{W}\|$ $<1.$ (8d)
@@ 3.2 A Proposed Choice of Lagrange Multipliers
Theorem 3.1 presents a specialization of the Karush-Kuhn-Tucker (KKT) optimality conditions for the semidefinite program (3) (see Section 5.5.3 of Boyd (2004)). Our derivation of sufficient conditions for perfect recovery relies on solving the nonlinear system (8), which ensures both optimality and uniqueness of the solution given by Theorem 3.1. This solution corresponds to the planted submatrix with rows and columns indexed by $U_{1}$ and $V_{1}$, respectively. Let $\boldsymbol{\bar{u}}$ and $\boldsymbol{\bar{v}}$ denote the characteristic vectors of $\bar{U}_{1}$ and $\bar{V}_{1}$.
Under the assumptions of Theorem 2.1, we show that the solutions $\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ are the unique optimal solutions of (3), and that the subgraph induced by the index sets $(m_{1},n_{1})$ represents the densest subgraph. Moreover, we demonstrate that $\{\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})\}$ satisfy the sufficient conditions for optimality derived from the KKT framework.
@@ 3.2.1 Choice of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$
    The complementary slackness condition (8c) is only satisfied in this case if $\Xi_{ij}=0.$ In this case,
@@
    $W_{ij}=\psi_{rs}\lambda\times\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega,\end{cases}$ (9)
    where
    $\psi_{rs}=\begin{cases}1,&\text{if }p_{rs}\leq 1-p_{rs}\\ -\frac{1-p_{rs}}{p_{rs}},&\text{if }1-p_{rs}<p_{rs}.\end{cases}$ (10)
    We choose $\Xi_{ij}=W_{ij}-\lambda+\gamma$ according to (8a).
  • Case 4: Suppose that $i\in U_{1}$, $j\in V-V_{1}$.
    Suppose further that $p_{1s}\leq 1/2.$ In this case, we choose so $W_{ij}$ and $\Xi_{ij}$ as follows so that (8a) is satisfied:
@@
    On the other hand, if $p_{1s}>1/2$, we choose
    $W_{ij}$ $=\begin{cases}\lambda\left(\frac{\nu_{j}-m_{1}}{\nu_{j}}\right),&\text{if }(i,j)\notin\Omega,\\ \lambda,&\text{if }(i,j)\in\Omega,\end{cases}$
    $\Xi_{ij}$ $=\begin{cases}\gamma-\lambda\left(\frac{m_{1}}{\nu_{j}}\right),&\text{if }(i,j)\notin\Omega,\\ \gamma,&\text{if }(i,j)\in\Omega.\end{cases}$
  • Case 5: Similarly, if $i\in U_{r}$, $r\geq 2$, $j\in V_{1}$ and $p_{r1}\leq 1/2$ we choose
@@
    If $p_{r1}>1/2$, we choose
    $W_{ij}$ $=\begin{cases}\lambda\left(\frac{\mu_{i}-n_{1}}{\mu_{i}}\right),&\text{if }(i,j)\notin\Omega,\\ \lambda,&\text{if }(i,j)\in\Omega.\end{cases}$
    $\Xi_{ij}$ $=\begin{cases}\gamma-\lambda\left(\frac{n_{1}}{\mu_{i}}\right),&\text{if }(i,j)\notin\Omega,\\ \gamma,&\text{if }(i,j)\in\Omega.\end{cases}$
@@ 3.2.2 Construction of $\boldsymbol{\Lambda}$
By construction, the proposed Lagrange multipliers satisfy the stationarity condition (8a) and the complementary slackness conditions (8b) and (8c). Moreover, for any $i\in U-U_{1}$, we have
@@ 3.4.1 Nonnegativity in Case 3
Consider $i\in U_{r}$ and $j\in V_{s}$ for $r,s\geq 2$. To determine sufficient conditions for nonnegativity of $\Xi_{ij}$, we must consider two cases dictating the choice of $\psi_{rs}$ and, hence, $W_{ij}$.
First, suppose that $p_{rs}\leq 1-p_{rs}$ or, equivalently, $p_{rs}\leq 1/2$. Recall that we choose $\psi=\psi_{rs}=1$ in this case. If $(i,j)\notin\Omega$, then
$\Xi_{ij}=W_{ij}-\lambda+\gamma=\lambda\psi-\lambda+\gamma=\gamma>0,$
by our choice of $W_{ij}$.
On the other hand, suppose that $(i,j)\in\Omega$. In this case,
$\Xi_{ij}$ $=W_{ij}-\lambda+\gamma=\gamma-\lambda\left(\psi\frac{p_{rs}}{1-p_{rs}}+1\right).$
@@
We next consider the case when $p_{rs}>1/2$. In this case,
$\psi=\psi_{rs}=-\frac{1-p_{rs}}{p_{rs}}.$
If $(i,j)\notin\Omega$, we have
$\Xi_{ij}=W_{ij}-\lambda+\gamma=-\frac{\lambda}{p_{rs}}+\gamma.$
It follows that
$p_{rs}\Xi_{ij}$ $=-\lambda+p_{rs}\gamma$
$=p_{rs}\gamma-\frac{1}{\sqrt{m_{1}n_{1}}}-\gamma(1-p_{11})-\gamma\tau$
$>(1-p_{rs})\gamma-\frac{1}{\sqrt{m_{1}n_{1}}}-\gamma(1-p_{11})-\gamma\tau$
$\geq\frac{\gamma}{2}(p_{11}-p_{rs})-\frac{1}{\sqrt{m_{1}n_{1}}},$
under Assumption (30) since $p_{rs}>1-p_{rs}$. Thus, $\Xi_{ij}>0$ under Assumptions (30) and (31) in this case. Finally, if $(i,j)\in\Omega$ then
$\Xi_{ij}=-\lambda\left(\frac{p_{rs}}{1-p_{rs}}\right)\left(-\frac{1-p_{rs}}{p_{rs}}\right)-\lambda+\gamma=\gamma>0.$
@@ 3.4.2 Nonnegativity in Cases 4 and 5
We start with Case 4 in the construction of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$. Fix $i\in U_{1}$ and $j\in V_{s}$ for $s\geq 2$.
We first consider the case when $p_{1s}>1/2$ and $(i,j)\notin\Omega$. In this case,
$\Xi_{ij}$ $=\gamma-\lambda\frac{m_{1}}{\nu_{j}}.$ (32)
Note that $\Xi_{ij}$ is a monotonically increasing function of $\nu_{j}$ on the interval $0,m_{j})$. Therefore, $\Xi_{ij}$ is minimized when $\nu_{j}$ is minimized over this interval. Moreover, the Bernstein inequality implies that
$\nu_{j}$ $\geq p_{1s}m_{1}-6\max\left\{\sqrt{\sigma_{p_{1s}}^{2}m_{1}\log{N^{*}}},\log{N^{*}}\right\}.$ (33)
with high probability. Substituting (29) into (32) and rearranging we see that $\Xi_{ij}\geq 0$ w.h.p. if and only if $(\nu_{j}/m_{1})\Xi_{ij}\geq 0$, which holds if
$\frac{\nu_{j}}{m_{1}}\Xi_{ij}$ $\geq\frac{\gamma}{m_{1}}\left(p_{1s}m_{1}-6\max\left\{\sqrt{\sigma_{1s}^{2}m_{1}\log N^{*}},\log N^{*}\right\}\right)-\gamma\tau-\gamma(1-p_{11})-\frac{1}{\sqrt{m_{1}n_{1}}}$
$\geq\gamma\left(p_{11}-p_{1s}-\tau-6\max\left\{\sqrt{\sigma_{1s}^{2}m_{1}\log N^{*}},\log N^{*}\right\}\right)-\frac{1}{\sqrt{m_{1}n_{1}}}.$
@@ 3.5.1 A Decomposition of $\boldsymbol{W}$
$\Pr\left(\lVert{\boldsymbol{A}}\rVert\geq C\max\left\{\sigma_{1}+\sigma_{2},\sqrt{B\log{t}}\right\}\right)\leq\max\{n_{r},m_{s}\}t^{-7}$
@@
Lemma 3.2 is a special case of the matrix concentration inequality given in (Van Handel and Youssef, 2017, Remark 4.11). Note there is no assumption of symmetry, i.e., that $a_{ij}=a_{ji}$, in the hypothesis of Lemma 3.2.
@@ 3.5.3 A Bound on $\boldsymbol{S}$
$\|\boldsymbol{S}\|\leq c_{S}\max\left\{\sqrt{\tilde{\sigma}^{2}_{*}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}$ (44)
@@
    $\boldsymbol{S}_{1}]_{ij}=S_{ij}=\frac{W_{ij}}{\lambda}=\psi_{rs}\times\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega,\end{cases}$
    where $\psi_{rs}$ is defined as in (10).
@@
    $\boldsymbol{S}_{1}]_{ij}=\psi_{1s}\times\begin{cases}1,&\text{with probability $p_{1s}$},\\ -\frac{p_{1s}}{1-p_{1s}},&\text{with probability }1-p_{1s},\end{cases}$
    where
    $\psi_{1s}=\begin{cases}1,&\text{if }p_{1s}\leq 1-p_{1s}\\ -\frac{1-p_{1s}}{p_{1s}},&\text{if }1-p_{1s}<p_{1s}.\end{cases}$
@@
    $\boldsymbol{S}_{1}]_{ij}=\psi_{r1}\times\begin{cases}1,&\text{with probability $p_{r1}$},\\ -\frac{p_{r1}}{1-p_{r1}},&\text{with probability }1-p_{r1},\end{cases}$
    where
    $\psi_{r1}=\begin{cases}1,&\text{if }p_{r1}\leq 1-p_{r1}\\ -\frac{1-p_{r1}}{p_{r1}},&\text{if }1-p_{r1}<p_{r1}.\end{cases}$
@@
$\tilde{\sigma}_{rs}^{2}:=\min\left\{\frac{p_{rs}}{1-p_{rs}},\frac{1-p_{rs}}{p_{rs}}\right\}$
@@
$\tilde{\sigma}^{2}=\max_{\begin{subarray}{c}r,s\geq 2\\ r\neq s\end{subarray}}\left\{\tilde{\sigma}^{2}_{rs}\right\}$
Applying Lemma 3.2 with $B=\max\{1,\tilde{\sigma}^{2}\}\leq 1$ and $t=N^{*}=\max\{N,M\}$, shows that
$\lVert\boldsymbol{S}_{1}\rVert=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}},\sqrt{\log N^{*}}\right\}\right)$ (45)
@@
for all $i\in[n]$, $j\in V_{s}$ for probability $p_{s}\in[0,1/2]$. Let $\tilde{\boldsymbol{\Theta}}$ be the random matrix defined by
@@
Note that $\boldsymbol{S}_{2}$ has the same form as $\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}$ in Lemma 3.5. Indeed, if $p_{1s}\leq 1/2$ then $S_{ij}$ is sampled according to the distribution defining $\boldsymbol{\Theta}$ given in (46) and $\boldsymbol{S}_{1}]_{ij}$ is sampled according to (47) with respect to $p_{s}=p_{1s}$. On the other hand, if $p_{1s}>1/2$ then $S_{ij}$ and $\boldsymbol{S}_{1}]_{ij}$ are sampled according to (46) and (47), respectively, with $p_{s}=1-p_{1s}<1/2$. Thus, $\boldsymbol{S}_{2}$ is of the form $\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}$ in Lemma 3.5 with $n=n_{1}$, $N=N-n_{1}$. It follows immediately that
@@
$\|\boldsymbol{S}\|=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}_{*}N^{*}\log N^{*}},(\log N^{*})^{3/2}\right\}\right)$

### v2

@@ 3.1 Optimality Conditions
$\Xi_{ij}$ $=\gamma\;\;\text{for all }(i,j)\notin\Omega$ (9b)
$\mathrm{Tr}(\boldsymbol{\Lambda}^{T}(\boldsymbol{\bar{X}}-\boldsymbol{1}\boldsymbol{1}^{T}))$ $=0$ (9c)
$\mathrm{Tr}(\boldsymbol{\Xi}^{T}\boldsymbol{\bar{Y}})$ $=0$ (9d)
$\boldsymbol{W}^{T}\boldsymbol{\bar{u}}=\boldsymbol{0},\quad\boldsymbol{W}\boldsymbol{\bar{v}}=\boldsymbol{0},\quad\|\boldsymbol{W}\|$ $<1.$ (9e)
Here, $\lambda$, $\boldsymbol{\Xi}$, $\boldsymbol{\Lambda}$ are Lagrange multipliers corresponding to the constraints $\mathrm{Tr}(\boldsymbol{X}\boldsymbol{1}\boldsymbol{1}^{T})=mn$, $\boldsymbol{Y}\geq 0$, and $\boldsymbol{X}\leq 1$, respectively, in (3); the remaining constraints in (3) and the corresponding Lagrange multipliers can be eliminated when formulating and applying the KKT conditions at $(\boldsymbol{\bar{u}},\boldsymbol{\bar{v}})$. Conditions (9a) and (9b) are the stationary conditions with respect to $\boldsymbol{X}$ and $\boldsymbol{Y}$. We note that the subdifferential of the nuclear norm at $\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ is given by
$\partial\|\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}\|_{*}=\left\{\frac{\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}}{\sqrt{m_{1}n_{1}}}+\boldsymbol{W}:\boldsymbol{W}^{T}\boldsymbol{\bar{u}}=\boldsymbol{0},\quad\boldsymbol{W}\boldsymbol{\bar{v}}=\boldsymbol{0},\quad\|\boldsymbol{W}\|\leq 1\right\},$
which corresponds to the first term in the stationarity condition (9a) and the inequalities (9e). We next consider a particular choice of these Lagrange multipliers and subgradient, and then determine conditions ensuring that the system of equations (9) is satisfied.
@@ 3.2 A Proposed Choice of Lagrange Multipliers
Theorem 3.1 presents a specialization of the Karush-Kuhn-Tucker (KKT) optimality conditions for the convex program (3) (see Section 5.5.3 of Boyd (2004)). Our derivation of sufficient conditions for perfect recovery relies on solving the nonlinear system (9), which ensures both optimality and uniqueness of the solution given by Theorem 3.1. This solution corresponds to the planted submatrix with rows and columns indexed by $U_{1}$ and $V_{1}$, respectively. Let $\boldsymbol{\bar{u}}$ and $\boldsymbol{\bar{v}}$ denote the characteristic vectors of $\bar{U}_{1}$ and $\bar{V}_{1}$.
Under the assumptions of Theorem 2.1, we show that the solutions $\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ are the unique optimal solutions of (3), and that the subgraph induced by the index sets $(m_{1},n_{1})$ represents the densest subgraph. Moreover, we demonstrate that $\bar{\boldsymbol{X}}=\boldsymbol{\bar{u}}\boldsymbol{\bar{v}}^{T}$ and $\bar{\boldsymbol{Y}}=P_{\Omega}(\bar{\boldsymbol{X}})$ satisfy the sufficient conditions for optimality derived from the KKT framework.
@@ 3.2.1 Choice of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$
    The complementary slackness condition (9d) is only satisfied in this case if $\Xi_{ij}=0.$ In this case,
@@
    $W_{ij}=\lambda\cdot\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega.\end{cases}$ (10)
    We choose $\Xi_{ij}=W_{ij}-\lambda+\gamma$ according to (9a) and (9b).
  • Case 4: Suppose that $i\in U_{1}$, $j\in V_{s}$ for $s\geq 2$. In this case, we choose $W_{ij}$ and $\Xi_{ij}$ as follows so that (9a) and (9b) are satisfied:
@@
  • Case 5: Similarly, if $i\in U_{r}$, $r\geq 2$, $j\in V_{1}$ we choose
@@ 3.2.2 Construction of $\boldsymbol{\Lambda}$
By construction, the proposed Lagrange multipliers satisfy the stationarity condition (9a) and the complementary slackness conditions (9c) and (9d). Moreover, for any $i\in U-U_{1}$, we have
@@ 3.4.1 Nonnegativity in Case 3
Consider $i\in U_{r}$ and $j\in V_{s}$ for $r,s\geq 2$.
If $(i,j)\notin\Omega$, then we must have $\Xi_{ij}=\gamma>0$ by (9b). On the other hand, suppose that $(i,j)\in\Omega$. In this case,
$\Xi_{ij}$ $=W_{ij}-\lambda+\gamma=\gamma-\lambda\left(\frac{p_{rs}}{1-p_{rs}}+1\right)=\gamma-\frac{\lambda}{1-p_{rs}}.$
@@ 3.4.2 Nonnegativity in Cases 4 and 5
We start with Case 4 in the construction of $\boldsymbol{W}$ and $\boldsymbol{\Xi}$. Fix $i\in U_{1}$ and $j\in V_{s}$ for $s\geq 2$. Then,
$\Xi_{ij}$ $=\gamma-\lambda\frac{m_{1}}{m-\nu_{j}}.$ (32)
Note that $\Xi_{ij}$ is a monotonically decreasing function of $\nu_{j}$ on the interval $0,m_{j})$. Therefore, $\Xi_{ij}$ is minimized when $\nu_{j}$ is maximized over this interval. Moreover, the Bernstein inequality implies that
$\nu_{j}$ $\leq p_{1s}m_{1}+6\max\left\{\sqrt{\sigma_{p_{1s}}^{2}m_{1}\log{N^{*}}},\log{N^{*}}\right\}.$ (33)
with high probability. Substituting (29) into (32) and rearranging we see that $\Xi_{ij}\geq 0$ w.h.p. if
$\gamma\left(p_{11}-p_{1s}-\tau-6\max\left\{\sqrt{\sigma_{1s}^{2}\frac{\log N^{*}}{m_{1}}},\frac{\log N^{*}}{m_{1}}\right\}\right)$ $\geq\frac{1}{\sqrt{m_{1}n_{1}}}.$
@@ 3.5.1 A Decomposition of $\boldsymbol{W}$
$\Pr\left(\lVert{\boldsymbol{A}}\rVert\geq C\max\left\{\sigma_{1}+\sigma_{2},\sqrt{B\log{t}}\right\}\right)\leq\max\{n,m\}t^{-7}$
@@
Lemma 3.2 is a special case of the matrix concentration inequality given in (Latała et al., 2018, Remark 4.12). Note there is no assumption of symmetry, i.e., that $a_{ij}=a_{ji}$, in the hypothesis of Lemma 3.2.
@@ 3.5.3 A Bound on $\boldsymbol{S}$
$\|\boldsymbol{S}\|\leq c_{S}\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}\log N^{*}},\sqrt{\max\{1,\tilde{\sigma}^{2}\}\log N^{*}},(\log N^{*})^{3/2}\right\}$ (44)
@@
    $\boldsymbol{S}_{1}]_{ij}=S_{ij}=\frac{W_{ij}}{\lambda}\begin{cases}1,&\text{if }(i,j)\notin\Omega,\\ -\frac{p_{rs}}{1-p_{rs}},&\text{if }(i,j)\in\Omega.\end{cases}$
@@
    $\boldsymbol{S}_{1}]_{ij}=\begin{cases}1,&\text{with probability $p_{1s}$},\\ -\frac{p_{1s}}{1-p_{1s}},&\text{with probability }1-p_{1s}.\end{cases}$
@@
    $\boldsymbol{S}_{1}]_{ij}=\psi_{r1}\times\begin{cases}1,&\text{with probability $p_{r1}$},\\ -\frac{p_{r1}}{1-p_{r1}},&\text{with probability }1-p_{r1}.\end{cases}$
@@
$\tilde{\sigma}_{rs}^{2}:=\frac{p_{rs}}{1-p_{rs}}$
@@
$\tilde{\sigma}^{2}=\max_{\begin{subarray}{c}(r,s)\neq(1,1)\end{subarray}}\left\{\tilde{\sigma}^{2}_{rs}\right\}$
Applying Lemma 3.2 with $B=\max\{1,\tilde{\sigma}^{2}\}$ and $t=N^{*}=\max\{N,M\}$, shows that
$\lVert\boldsymbol{S}_{1}\rVert=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}},\sqrt{B\log N^{*}}\right\}\right)$ (45)
@@
for all $i\in[n]$, $j\in V_{s}$ for probability $p_{s}\in[0,1]$. Let $\tilde{\boldsymbol{\Theta}}$ be the random matrix defined by
@@
Note that $\boldsymbol{S}_{2}$ has the same form as $\boldsymbol{\Theta}-\tilde{\boldsymbol{\Theta}}$ in Lemma 3.5. Indeed, $S_{ij}$ is sampled according to the distribution defining $\boldsymbol{\Theta}$ given in (46) and $\boldsymbol{S}_{1}]_{ij}$ is sampled according to (47) with respect to $p_{s}=p_{1s}$ for each $j\in V_{s}$. It follows immediately that
@@
$\|\boldsymbol{S}\|=\mathcal{O}\left(\max\left\{\sqrt{\tilde{\sigma}^{2}N^{*}\log N^{*}},\sqrt{\max\{1,\tilde{\sigma}^{2}\}\log N^{*}},(\log N^{*})^{3/2}\right\}\right)$